How many zeros are there in one crore?
A. 5
B. 6
C. 7
D. 8

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
One crore can also be called as ten million and the hundred lakhs. Here we will represent the one crore in the mathematical form. Here we will also take the help of the thousands and lakhs terms to get the required value. Complete step-by-step answer: Since, we don’t know the number of zeros in crores – first we will take the help of thousands. Ten hundred makes one thousand. 10×100=1000 Hundred thousand makes one lakh. 100×1000=1,00,000 Ten lakhs make one million. 1,00,000×10=1,000,000 Ten million makes one crore. 10×1,000,000=1,00,00,000 Hence, we can count the number of zeros in the number of crore and we can see that crore has seven zeros. Hence, from the given multiple choices – the option C is the correct answer. So, the correct answer is “Option C”. Note: Number system can be defined as the writing system for expressing numbers and represents the numbers for the given set, using digits or the other symbols in the consistent manner. There are two types of number systems. They are Indian number system and the international number system. It gives the unique representation of every number.